Ruby has just finished her economics exam and must decide whether to
spend two hours going to the beauty parlour, having a couple of pizzas, or
going to a movie. She can do one, all, or any two of these things. The question
is, what should she do? The marginal utility of the beauty parlour service is
36 utils and the price is $6. The marginal utility of the two pizzas is 24 utils
and the price is $4 for two pizzas. The marginal utility of going to the movie
is 40 utils and the price is $10. Ruby's marginal utility of a dollar is 5 utils per
dollar.
1
Expert's answer
2021-09-14T18:57:29-0400

Given the marginal utility of beauty parlor = 36 utils

Marginal utility from pizza = 24 utils

Marginal utility from movie = 40 utils

"Per\\space dollar\\space MU\u2009\\space in\\space the\\space case\\space of\\space Beauty\\space parlour = \\frac{36}{6}=6 \\space utils\\\\\n\nPer\\space dollar\\space MU\u2009\\space in\\space the\\space case\\space of\\space pizza = \\frac{24}{4}=6 \\space utils\\\\\n\nPer\\space dollar\\space MU\u2009\\space in\\space the\\space case\\space of\\space movie = \\frac{40}{10}=4 \\space utils"

She will go for Pizza and beauty parlor because the MU per dollar from these two is greater than Ruby's marginal utility of a dollar.